what is the molarity of a HCl solution that contains 6.40g of HCl in 477mL of solution? answer in units of M
1
Expert's answer
2016-04-18T10:04:05-0400
M(HCl) = 35.5 + 1 = 36.5 (g/mol)
n(HCl) = m/M = 6.4 / 36.5 = 0.175 (mol)
c(HCl) = n / c = 0.175 / 0.477 = 0.37 (mol/l)
Answer
c(HCl) = 0.37 M